# Dynamic Contrastive Knowledge Distillation for Efficient Image Restoration

### v1

@@ Conclusion
In this work, we propose a dynamic contrastive knowledge distillation framework for image restoration, named DCKD, which consists of Dynamic Contrastive Regularization (DCR) and Distribution Mapping Module (DMM). Most previous knowledge distillation methods utilize a fixed solution space, causing the lower bound constraints to weaken gradually during training. DCR constructs a dynamic solution space based on the student’s learning state to enhance lower-bound constraints. DMM is the first time introducing pixel-level category information for knowledge distillation in low-level vision tasks. Experiments on image super-resolution, image deraining, and image deblurring tasks validate that the proposed DCKD achieves state-of-the-art results on various benchmark datasets both quantitatively and visually.

### v2

@@ Conclusion
In this work, we propose a dynamic contrastive knowledge distillation framework for image restoration, named DCKD, which consists of the Dynamic Contrastive Regularization (DCR) and the Distribution Mapping Module (DMM). Most previous knowledge distillation methods utilize a fixed solution space, causing the lower bound constraints to weaken gradually during training. DCR constructs a dynamic solution space based on the student’s learning state to enhance the lower-bound constraints. DMM introduces pixel-level category information to knowledge distillation for low-level vision tasks for the first time. Experiments on image super-resolution, image deblurring, and image deraining tasks validate that the proposed DCKD achieves state-of-the-art results on various benchmark datasets, both quantitatively and visually.
Model Scale Method Set14 BSD100 Urban100 Manga109
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
SwinIR-light $\times$4 - 28.77/0.7858 27.69/0.7406 26.47/0.7980 30.92/0.9151
MCLIR 28.85/0.7874 27.72/0.7414 26.57/0.8010 31.04/0.9158
DCKD 28.88/0.7891 27.75/0.7432 26.64/0.8039 31.20/0.9180
Table 8: Quantitative comparison DCKD with MCLIR (Wu et al. 2024) on the benchmark datasets for SR task.
Figure 9: More visual comparison of image super-resolution at ×$\times 4$.
Figure 10: More visual comparison of image deblurring.
Figure 11: More visual comparison of image deraining.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (NO. 62102151), the Open Research Fund of Key Laboratory of Advanced Theory and Application in Statistics and Data Science, Ministry of Education (KLATASDS2305), the Fundamental Research Funds for the Central Universities.
Appendix
Implementation Details
The details of image encoder.
For the latent features, we extract the features from five different layers of the pre-trained VQGAN (Esser, Rombach, and Ommer 2021), while with the corresponding coefficients $\lambda_{i},i=1,\cdots 5$ to $\frac{1}{32},\frac{1}{16},\frac{1}{8},\frac{1}{4}$ and 1, respectively. We set the attenuation rate $\alpha$ to 0.1.
The details of different backbones.
For NAFNet (Chen et al. 2022a), we train the model with AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.9$, weight deacy = $10^{-3}$) for total $2\times 10^{5}$ iterations with the initial learning rate $10^{-3}$ gradually reduced to $10^{-7}$ with the cosine annealing schedule (Loshchilov and Hutter 2016). The training patch size is 256 × 256 and the batch size is 32.
For Restormer (Zamir et al. 2022), we train the model with AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$, weight deacy = $10^{-4}$) for $3\times 10^{5}$ iterations with the initial learning rate $3\times 10^{-4}$ gradually reduced to $1\times 10^{-6}$ with the cosine annealing (Loshchilov and Hutter 2016). We start training with patch size $128\times 128$ and batch size 64 for progressive learning. The patch size and batch size pairs are updated to [$(160^{2},40),(192^{2},32),(256^{2},16),(320^{2},8),(384^{2},8)$] at iterations [92K, 156K, 204K, 240K, 276K]. For data augmentation, we use horizontal and vertical flips.
Comparison of DCKD with Contrastive Learning
The quantitative comparison of DCKD with contrastive learning on the benchmark datasets for the SR task is reported in Tab. 8. DCKD significantly surpasses state-of-the-art method MCLIR (Wu et al. 2024) 0.16dB on Manga109. This shows that DCKD provides more comprehensive lower-bound constraints, ensuring that each negative sample exerts a consistent force on the anchor sample. Additionally, our method requires only a single negative sample model to generate an arbitrary number of negative samples, significantly reducing computational overhead.
More Visual Comparison
We provide more visual comparisons. Fig. 9 presents some challenging images from the x4 super-resolution task. As observed, previous distillation methods often result in blurred artifacts and a struggle to effectively restore high-frequency details such as lines and architectural structures. For example, in img002 and img026, previous methods produce significant artifacts between lines or blur the gaps between them. In contrast, our DCKD not only removes these artifacts but also clearly separates the lines. Overall, our method enhances the student model’s ability to handle high-frequency details and accurately restore them.
Further visual comparisons for deblurring and deraining tasks are shown in Fig. 10 and Fig. 11, where our DCKD effectively removes motion blur and rain, with the restored image quality closely matching that of the ground truth (GT).